# Accelerating Diffusion Transformers with Dual Feature Caching

### v1

@@
Abstract
Diffusion Transformers (DiT) have become the dominant methods in image and video generation yet still suffer substantial computational costs. As an effective approach for DiT acceleration, feature caching methods are designed to cache the features of DiT in previous timesteps and reuse them in the next timesteps, allowing us to skip the computation in the next timesteps. However, on the one hand, aggressively reusing all the features cached in previous timesteps leads to a severe drop in generation quality. On the other hand, conservatively caching only the features in the redundant layers or tokens but still computing the important ones successfully preserves the generation quality but results in reductions in acceleration ratios. Observing such a tradeoff between generation quality and acceleration performance, this paper begins by quantitatively studying the accumulated error from cached features. Surprisingly, we find that aggressive caching does not introduce significantly more caching errors in the caching step, and the conservative feature caching can fix the error introduced by aggressive caching. Thereby, we propose a dual caching strategy that adopts aggressive and conservative caching iteratively, leading to significant acceleration and high generation quality at the same time. Besides, we further introduce a V-caching strategy for token-wise conservative caching, which is compatible with flash attention and requires no training and calibration data.
Our codes have been released in Github:
Code: https://github.com/Shenyi-Z/DuCa
1 Introduction
@@
Diffusion Models (DMs)[16] have shown remarkable success in both image [27] and video generation [2]. Traditional DMs often use U-Net denoising structures, such as Stable Diffusion (SD)[27] and Stable Video Diffusion (SVD)[2], which have already demonstrated impressive results. Most recently, with the introduction of Diffusion Transformers (DiT)[26], the quality of visual generation has further significantly improved, propelling the development of numerous downstream applications [12]. However, the quality improvement is also accompanied by a significant increase in computational demands, reducing the inference efficiency of transformers and making their deployment in practical scenarios a pressing challenge.
@@
Recently, feature caching has become one of the most popular techniques for diffusion acceleration. Motivated by the high similarity between features in the adjacent timesteps, aggressive feature caching caches the features computed in a freshing timestep, and then reuses them in the following several caching steps, as shown in Figure 1(b). Such a fresh-then-cache strategy enables the diffusion model to save computational costs in the caching step and thus leads to significant acceleration. However, since the similarity of features in non-adjacent timesteps decreases quickly, the error introduced by aggressive feature caching can be accumulated at an exponential speed, leading to a significant drop in generation quality, and making aggressive feature caching impractical.
@@
Figure 1: Comparison between conservative caching (a) and aggressive caching (b): Conservative caching skips fewer layers than aggressive caching to maintain the generation quality, but also leads to lower acceleration ratios. $t$ and $t-1$ denote the refreshing and caching timesteps, respectively.
To solve this problem, recent works [8, 24, 28, 35] introduce a few computations during the caching steps in certain layers and tokens, which make the cached features in previous steps be corrected for the current caching step. These conservative feature caching methods reduce the accumulated error from feature caching but also make it difficult to achieve a higher acceleration ratio. Such a tradeoff between acceleration ratio and generation quality becomes the most essential problem for feature caching methods.
@@
To investigate this issue, we first examine the performance of conservative caching methods and the aggressive caching method quantitatively. As shown in Figure 2, we observe that the caching error from conservative caching methods increases at a relatively slower rate as the caching step grows, while the error in the aggressive caching method accumulates rapidly with each additional caching step, which explains the aforementioned tradeoff. Notably, we find that the caching error of the aggressive and conservative caching exhibit a close value at the first caching step (i.e., timestep 27 in Figure 2) but a significantly larger value in the following caching steps, suggesting that it is possible to have high acceleration ratios and generation quality by leveraging both cache schemes at appropriate caching steps.
Based on this observation, we propose DUal CAching, which leverages the advantages of aggressive and conservative caching methods. In each caching cycle, DuCa first initializes the cache by performing the full computation in the fresh timestep. Then, it performs the one-step aggressive caching for high-ratio acceleration, followed by one-step conservative caching to fix the cache error. Such an alternative caching strategy can be repeated multiple times.
In addition, we further introduce a token-wise caching method as our conservative caching strategy. Token-wise feature caching has been recently proposed to perform computation for only the important tokens in the caching steps. However, the definition of important tokens in the previous work [35] relies on the attention weights in self-attention, making it not compatible with FlashAttention[9] and memory-efficient attention. This problem further improves the memory costs of attention from $O(N)$ to $O(N^{2})$, also resulting in additional latency, making it not applicable. To address this issue, we introduce V-Caching, which approximates the attention weight-based importance score by the norm of the value matrix, providing good compatibility with FlashAttention and memory-efficient Attention while maintaining good generation quality.
Extensive experiments have demonstrated the performance of our method in abundant benchmarks, including text-to-image generation with PixArt-$\alpha$, text-to-video generation with OpenSora, and class-conditioned image generation with DiT. For instance, 1.98×1.98$\times$1.98 × acceleration on PixArt-$\alpha$ and 2.50×2.50$\times$2.50 × acceleration on OpenSora with almost no loss on generation quality.
In summary, our main contributions are as follows:
@@
    We study the caching error for aggressive feature caching and conservative feature caching, and introduce DuCa, a dual feature caching strategy to perform the two caching methods alternatively, maintaining the generation quality while achieving a higher acceleration ratio.
@@
    We introduce V-caching as the conservative caching strategy in DuCa, a token-wise feature caching strategy that selects the important tokens via the norm of their value matrix, making it compatible with FlashAttention[9].
@@
2 Related Works
Transformers in Diffusion Models Diffusion Models (DMs)[16, 30] generate images by gradually adding noise to the image during training and iteratively applying a denoising model to a randomly initialized noise during inference. DMs have recently gained extensive application across various tasks in image generation due to their outstanding performance [27, 1]. Earlier DMs with U-Net-based architectures [16, 27], which also include some Transformer layers, have achieved impressive results. Recently, Diffusion Transformer (DiT) [26] has been constructed by directly stacking multiple DiT Blocks in a cascade, achieving performance in class-conditional image generation that surpasses U-Net-based DMs. Consequently, the DiT architecture has been further extended to several generative fields, yielding exceptional outcomes. For example, the PixArt series [7, 6, 5] and similar text-to-image models incorporate Cross-Attention into the DiT Block to add text conditional information into generation procedure, while the Sora series [4, 18] enhances the DiT Block’s Attention structure to Spatial-Temporal attention, enabling high-quality video generation. However, though success, these DiT-based DMs also entail significantly higher computational costs and time consumption, making deploying such generative models in real-world applications challenging.
@@ 2 Related Works
Acceleration of Diffusion Models To address the high computational costs commonly associated with DMs including DiT, various acceleration methods have been proposed. Existing acceleration methods can generally be divided into two categories: Designing sampling methods to reduce the number of steps needed to transition from noise to high-quality images, and directly accelerating the denoising model. Sampling-based acceleration methods enhance the efficiency of DMs by simply reducing the sampling steps. DDIM [31] leverages a deterministic, non-Markovian reverse process to significantly reduce generation steps while maintaining high-quality outputs. DPM Solver [22], through analytic derivation, optimizes the reverse process, enabling efficient sampling with fewer steps, while DPM Solver++ [23] further refines this approach by incorporating noise control and diverse sampling techniques to improve quality and adaptability. Rectified Flow [21] employs a recurrent flow framework, utilizing specially designed flow fields to accelerate the diffusion process.
@@
Feature Caching Methods that improve generation efficiency by accelerating the denoising model largely vary depending on the model structure. Acceleration methods for denoising models in different aspects have been proposed and work well, including weight quantization [29], structural pruning [11], knowledge distillation [19] and token merge [3]. Due to the high computational cost of retraining the diffusion models, or even small learnable modules introduce significant costs, limiting model reusability across different scales, resolutions, and tasks, driving interest in training-free acceleration methods. Faster Diffusion [27] improves computational efficiency by caching the output of the encoder module on certain steps, while DeepCache [25] reduces redundant computations by reusing low-resolution feature information in the skip connections of the U-Net structure. Unfortunately, although both are training-free methods, they are specialized feature caching techniques designed specifically for U-Net-based models like Stable Diffusion [27], making them difficult to directly adapt to transformer-based models like DiT [8]. Recently, to address the lack of feature reuse solutions for DiT, $\Delta$-DiT [8] and Pyramid Attention Broadcast (PAB) [33] have been introduced. The former focuses on constructing residuals of DiT outputs across different layers, while the latter emphasizes the importance of various types of attention within DiT Blocks containing multiple attention heads. Learning-to-Cache [24] achieves higher acceleration ratios by employing a learnable router to determine whether computations are needed at each layer, but it incurs prohibitive computational costs. These feature caching methods generally lack attention to finer-grained features. ToCa [35] proposes a method that selects important tokens at both layer and token granularity, allocating more computation to these tokens while caching less important ones, advancing temporal redundancy management to the token level.
The errors introduced by feature caching are primarily attributed to discrepancies in timestep-sensitive information. Comparing the feature caching methods for DiT discussed above, we find that these feature caching approaches for DiT employ different types of feature cache-reuse structures, resulting in significant variations in acceleration ratios while maintaining generation quality. Additionally, they commonly introduce additional computations to mitigate and correct the discrepancies between cached features and the requirements of the current time step. Our proposed method DuCa demonstrates that conservative architectures like ToCa, with their long-range feature reuse capability, can correct timestep information for aggressive methods that struggle with larger timestep intervals. By leveraging the high caching ratios of aggressive methods, DuCa further enhances acceleration efficiency.
3 Method
3.1 Preliminary
@@
Diffusion models [16] are designed with two main processes: a forward process, where Gaussian noise is incrementally added to a clean image, and a reverse process, where a standard Gaussian noise is gradually denoised to reconstruct the original image. Letting $t$ represent the timestep and $\beta_{t}$ the noise variance schedule, the conditional probability in the reverse (denoising) process, $p_{\theta}(x_{t-1}\mid x_{t})$, can be modeled as:
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{T}\alpha_{i}$, and $T$ represents the total number of timesteps. Notably, $\epsilon_{\theta}$ is a denoising network parameterized by $\theta$, which takes $x_{t}$ and $t$ as the inputs to predict the noise needed for denoising. In an image generation procedure using $\mathcal{T}$ timesteps, $\epsilon_{\theta}$ must perform inference $\mathcal{T}$ times, which constitutes the majority of computational costs in diffusion models. Recently, numerous studies have shown that implementing $\epsilon_{\theta}$ as a transformer often enhances generation quality.
Figure 3: The overview of DuCa. (a) In the first timestep of a cache cycle, DuCa computes all tokens and stores them in the cache as initialization. Then, DuCa performs a conservative caching step followed by an aggressive caching step, and then another conservative caching step, alternating in this manner to form a complete cache cycle. (b) The proposed V-Caching method selects important tokens for DuCa by first fetching the value matrix for all input tokens that have been previously computed in the attention layers and then computing the norm of the value matrix for each token. The tokens with the largest V-norm values are selected as cached tokens. The computation results for these tokens are directly retrieved from the cache, while the remaining tokens are computed through the neural layer, with their updated results stored back into the cache.
@@
$\mathcal{G}$, $g_{l}$, and $\mathcal{F}^{l}$ represent the DiT model, the blocks in DiT, and different layers in a DiT block, respectively. $l$ denotes the index of DiT blocks and $L$ denotes the depth of DiT. In diffusion transformers, the input data $\mathbf{x}_{t}$ consists of a sequence of tokens representing different patches in the generated images, expressed as $\mathbf{x}_{t}=\{x_{i}\}_{i=1}^{H\times W}$, where $H$ and $W$ indicate the height and width of the images or their latent code, respectively. Then, the computation between $\mathcal{F}$ and $\mathbf{x}$ can be generally written as $\mathcal{F}(\mathbf{x})=\mathbf{x}+\text{AdaLN}\circ f(\mathbf{x})$, where the first item $\mathbf{x}$ indicates the residual connection and $f$ denotes the MLP or attention layers. AdaLN indicates the adaptive layer normalization [13].
Feature Caching: Feature caching aims to reuse the features that are computed and cached in the past timesteps to skip the computation in future timesteps. It is usually performed in multiple repeated caching cycles. Assuming a caching cycle with $N$ timesteps from the timestep $t$ to $t+N-1$, then feature caching performs the full computation in the timestep $t$ and stores the obtained features in a cache. This progress can be denoted as $\mathcal{C}[l]:=g_{l}(x_{t})$, where $\mathcal{C}$ denotes the cache and “:=” denotes assigning value operation. Then, in timesteps from $t+1$ to $t+N-1$, the computation in these steps is skipped by reusing the cached features, which can be formulated as $g_{l}(x_{t+i}):=\mathcal{C}[l],i\leq N-1$. For simplicity, in this paper, the timestep $t$ and the timesteps in $[t+1,t+N-1]$ are named the freshing step and the caching steps, respectively.
3.2 Dual Feature Caching

### v2

# Rethinking Token-wise Feature Caching: Accelerating Diffusion Transformers with Dual Feature Caching

@@
Rethinking Token-wise Feature Caching:
Accelerating Diffusion Transformers with Dual Feature Caching
Chang Zou1,*, Shikang Zheng1,2,*, Evelyn Zhang1, Runlin Guo3, Haohang Xu4, Zhengyi Shi1
Conghui He5
, Xuming Hu6, Linfeng Zhang1,†
1
Shanghai Jiao Tong University 2 South China University of Technology 3 Beihang University
4 Huawei Technologies Ltd 5 Shanghai AI Lab 6 Hong Kong University of Science and Technology (Guangzhou)
* Equal contribution.
Denotes the corresponding author.
I Abstract
Diffusion Transformers (DiT) have become the dominant methods in image and video generation yet still suffer substantial computational costs. As an effective approach for DiT acceleration, feature caching methods are designed to cache the features of DiT in previous timesteps and reuse them in the next timesteps, allowing us to skip the computation in the next timesteps. Among them, token-wise feature caching has been introduced to perform different caching ratios for different tokens in DiTs, aiming to skip the computation for unimportant tokens while still computing the important ones. In this paper, we propose to carefully check the effectiveness in token-wise feature caching with the following two questions: (1) Is it really necessary to compute the so-called ”important” tokens in each step? (2) Are so-called important tokens really important? Surprisingly, this paper gives some counter-intuition answers, demonstrating that consistently computing the selected “important tokens” in all steps is not necessary. The selection of the so-called “important tokens” is often ineffective, and even sometimes shows inferior performance than random selection. Based on these observations, this paper introduces dual feature caching referred to as DuCa, which performs aggressive caching strategy and conservative caching strategy iteratively and selects the tokens for computing randomly. Extensive experimental results demonstrate the effectiveness of our method in DiT, PixArt, FLUX, and OpenSora, demonstrating significant improvements than the previous token-wise feature caching.
II Introduction
@@
Diffusion Models (DMs)[ho2020DDPM] have shown remarkable success in both image [rombach2022SD] and video generation [blattmann2023SVD]. Traditional DMs often use U-Net denoising structures, such as Stable Diffusion (SD)[rombach2022SD] and Stable Video Diffusion (SVD)[blattmann2023SVD], which have already demonstrated impressive results. Most recently, with the introduction of Diffusion Transformers (DiT)[peebles2023dit], the quality of visual generation has significantly improved further, propelling the development of numerous downstream applications [Feng2024DiT4EditDT]. However, the quality improvement is also accompanied by a significant increase in computational demands, reducing the inference efficiency of transformers and making their deployment in practical scenarios a pressing challenge.
@@
Recently, feature caching has become one of the most popular techniques for diffusion acceleration. Motivated by the high similarity between features in the adjacent timesteps, aggressive feature caching caches the features computed in a freshing timestep, and then reuse them in the following several caching steps, as shown in Figure 1(b). Such a fresh-then-cache strategy enables the diffusion model to save computational costs in the caching step and thus leads to significant acceleration. However, since the similarity of features in non-adjacent timesteps decreases quickly, the error introduced by aggressive feature caching can be accumulated at an exponential speed, leading to a significant drop in generation quality, and making aggressive feature caching impractical.
@@
Figure 1: Comparison between conservative caching (a) and aggressive caching (b): Conservative caching skips some tokens instead of all tokens in each layer to maintain the generation quality, but also leads to lower acceleration ratios. $t$ and t1t-1 denote the refreshing and caching timesteps, respectively.
To solve this problem, recent work [chen2024delta-dit, ma2024l2c, selvaraju2024fora, zou2024accelerating] introduces some computations during the caching steps in certain layers and tokens, which makes the cached features in previous steps corrected for the current caching step. Among them, token-wise feature caching (ToCa) has been proposed to selectively compute the important tokens while skipping the unimportant tokens with the cache, reducing the accumulated error from feature caching, but also making it difficult to achieve a higher acceleration ratio. In this paper, we rethink and distill ToCa. Specifically, we focus on the following two questions.
(I) Is it necessary to compute the important tokens in all caching steps? ToCa computes the important tokens in each caching step to reduce the accumulated error from feature caching. Figure 2 shows the caching error from the aggressive (marked in green) and conservative caching (ToCa, marked in blue). It is observed that the caching error from ToCa increases slowly as the caching step grows, while the error in the aggressive caching method accumulates rapidly with each caching step, which verifies the effectiveness of ToCa. However, we find that the caching error of the aggressive and conservative caching exhibit a close value at the first caching step (i.e., timestep 27 in Figure 2), suggesting that computing the important tokens in ToCa is not always necessary. Concretely, when the error from caching has not been accumulated yet, computing important tokens does not bring benefits and can be removed for better efficiency.
Solution: Dual Feature Caching: Based on this observation, we propose DUal CAching, which leverages the advantages of aggressive and conservative caching methods. In each caching cycle, DuCa initializes the cache by performing the full computation in the fresh timestep. Then, it performs the one-step aggressive caching for high-ratio acceleration, followed by one-step conservative caching to fix the cache error. This alternative caching strategy can be repeated multiple times to achieve the best balance between efficiency and quality.
(II) Are so-called important tokens really important? The basic assumption of ToCa is that there exist some tokens that are sensitive or robust to feature caching. Based on this motivation, ToCa introduces four kinds of scores to describe the importance of each token based on their attention scores, spatial position, and frequency of being cached, respectively. Unfortunately, these scores not only introduce additional computation costs, but also lead to the incompatibility of FlashAttention, one of the default efficient operators in hardware. Hence, it results in around 100% increments in latency and improves the memory overhead from $O(N)$ to $O(N^{2})$. To solve this problem, we carefully study the effectiveness of each score in ToCa. Surprisingly, as shown in Table I and will be discussed later, we find that selecting tokens with the highest attention scores in ToCa sometimes shows inferior performance than random token selection, indicating previous methods may misunderstand the importance of tokens.
Solution: Selecting tokens with low similarity: After carefully trying all kinds of importance scores for token selection, we find that choosing a subset of tokens that has the smallest similarity to other tokens is the only solution to beat random token selection. On the other hand, choosing a subset of tokens with maximal similarity to each other almost achieves the worst performance that we can have. Based on this observation, we argue that token selection should focus on their duplication instead of their importance. Thus, random selection can be a good choice since randomness guarantees choosing tokens with different semantic information and leads to no additional computation or memory overhead.
@@
Extensive experiments have demonstrated the performance of our method in abundant benchmarks, including text-to-image generation with FLUX.1-dev, PixArt-$\alpha$, text-to-video generation with OpenSora, and class-conditioned image generation with DiT. For instance, $3.45\times$ acceleration on FLUX.1-dev and $2.50\times$ acceleration on OpenSora with almost lossless quality. In summary, our main contributions are as follows:
@@
    We rethink the necessity of computing “important” tokens in token-wise feature caching, and introduce DuCa, a dual feature caching strategy to perform the two caching methods alternatively, maintaining the generation quality while achieving a higher acceleration ratio.
@@
    We rethink the tokens selection methods in token-wise feature caching, finding that previous selection methods show inferior performance than random token selection, and explaining this phenomenon from the perspective of duplication. By just replacing previous methods with random selection, we achieve good compatibility with FlashAttention, leading to $\sim 2\times$ speed up in latency.
@@
III Related Works
@@
Acceleration of Diffusion Models To address the high computational costs commonly associated with DMs including DiT, various acceleration methods have been proposed. Existing acceleration methods can generally be divided into two categories: Designing sampling methods to reduce the number of steps needed to transition from noise to high-quality images, and directly accelerating the denoising model. Sampling-based acceleration methods enhance the efficiency of DMs by simply reducing the sampling steps. DDIM [songDDIM] leverages a deterministic, non-Markovian reverse process to significantly reduce generation steps while maintaining high-quality outputs. DPM Solver [lu2022dpm], through analytic derivation, optimizes the reverse process, enabling efficient sampling with fewer steps, while DPM Solver++ [lu2022dpm++] further refines this approach by incorporating noise control and diverse sampling techniques to improve quality and adaptability. Rectified Flow [refitiedflow] employs a recurrent flow framework, utilizing specially designed flow fields to accelerate the diffusion process.
@@
Feature Caching Due to the high computational cost of retraining the diffusion models, or even small learnable modules introduce significant costs, limiting model reusability across different scales, resolutions, and tasks, driving interest in training-free acceleration methods. Faster Diffusion [rombach2022SD] improves computational efficiency by caching the output of the encoder module on certain steps, while DeepCache [ma2024deepcache] reduces redundant computations by reusing low-resolution feature information in the skip connections of the U-Net structure. Unfortunately, although both are training-free methods, they are specialized feature caching techniques designed specifically for U-Net-based models like Stable Diffusion [rombach2022SD], making them difficult to directly adapt to transformer-based models like DiT [chen2024delta-dit]. Recently, to address the lack of feature reuse solutions for DiT, $\Delta$-DiT [chen2024delta-dit] and Pyramid Attention Broadcast (PAB) [zhao2024PAB] have been introduced. The former focuses on constructing residuals of DiT outputs across different layers, while the latter emphasizes the importance of various types of attention within DiT Blocks containing multiple attention heads. Learning-to-Cache [ma2024l2c] achieves higher acceleration ratios by employing a learnable router to determine whether computations are needed at each layer, but it incurs prohibitive computational costs. These feature caching methods generally lack attention to finer-grained features. ToCa [zou2024accelerating] proposes a method that selects important tokens at both layer and token granularity, allocating more computation to these tokens while caching less important ones, advancing temporal redundancy management to the token level.
IV Method
IV-A Preliminary
@@
Diffusion models [ho2020DDPM] are designed with two main processes: a forward process, where Gaussian noise is incrementally added to a clean image, and a reverse process, where a standard Gaussian noise is gradually denoised to reconstruct the original image. Letting $t$ represent the timestep and $\beta_{t}$ the noise variance schedule, the conditional probability in the reverse (denoising) process, $p_{\theta}(x_{t-1}\mid x_{t})$, can be modeled as:
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{T}\alpha_{i}$, and $T$ represents the total number of timesteps. Notably, $\epsilon_{\theta}$ is a denoising network parameterized by $\theta$, which takes $x_{t}$ and $t$ as inputs to predict the noise needed for denoising. In an image generation procedure using $\mathcal{T}$ timesteps, $\epsilon_{\theta}$ must perform inference $\mathcal{T}$ times, which constitutes the majority of computational costs in diffusion models. Recent studies have shown that implementing $\epsilon_{\theta}$ as a transformer often improves the quality of generation.
Figure 3: The overview of DuCa. (a) In the first timestep of a cache cycle, DuCa computes all tokens and stores them in the cache as initialization. Then, DuCa performs a conservative caching step followed by an aggressive caching step, and then another conservative caching step, alternating in this manner to form a complete cache cycle. (b) The proposed method selects tokens for DuCa by setting random scores for all tokens to diversity. The tokens with the largest random scores are selected as cached tokens. The computation results for these tokens are directly retrieved from the cache, while the remaining tokens are computed through the neural layer, with their updated results stored back into the cache.
@@
$\mathcal{G}$, $g_{l}$, and $\mathcal{F}^{l}$ represent the DiT model, the blocks in DiT, and different layers in a DiT block, respectively. $l$ denotes the index of DiT blocks, and $L$ denotes the depth of DiT. In diffusion transformers, the input data $\mathbf{x}_{t}$ consists of a sequence of tokens representing different patches in the generated images, expressed as $\mathbf{x}_{t}=\{x_{i}\}_{i=1}^{H\times W}$, where $H$ and $W$ indicate the height and width of the images or their latent code, respectively. Then, the computation between $\mathcal{F}$ and $\mathbf{x}$ can be generally written as $\mathcal{F}(\mathbf{x})=\mathbf{x}+\text{AdaLN}\circ f(\mathbf{x})$, where the first item $\mathbf{x}$ indicates the residual connection and $f$ denotes the MLP or attention layers. AdaLN indicates the adaptive layer normalization [Guo2022AdaLNAV].
Feature Caching: Feature caching aims to reuse the features that are computed and cached in the past timesteps to skip the computation in future timesteps. It is usually performed in multiple repeated caching cycles. Assuming a caching cycle with $N$ timesteps from the timestep $t$ to t+N1t+N-1, then feature caching performs the full computation in the timestep $t$ and stores the obtained features in a cache. This progress can be denoted as $\mathcal{C}[l]:=g_{l}(x_{t})$, where $\mathcal{C}$ denotes the cache and “:=” denotes assigning value operation. Then, in timesteps from $t+1$ to t+N1t+N-1, the computation in these steps is skipped by reusing the cached features, which can be formulated as $g_{l}(x_{t+i}):=\mathcal{C}[l],i\leq N-1$. For simplicity, in this paper, the timestep $t$ and the timesteps in [t+1,t+N1][t+1,t+N-1] are named the freshing step and the caching steps, respectively.
IV-B Dual Feature Caching